Sales of tablet computers at Ted​ Glickman's electronics store in​ Washington, D.C., over the past 10 weeks are shown in thetable below:                                                   

                                                   Week

1 2 3 4 5 6 7 8 9 10

Demand

19 20 29 38 26 28 37 24 25 30 ​

Forecast

19.0

a) The forecast for weeks 2 through 10 using exponential smoothing with alpha ​= 0.50 and a week 1 initial forecast of 19.0 are ​(round your responses to two decimal​ places):
Answer #1
Week Demand Forecast
1 19 19.00
2 20 =(0.5*19+0.5*19) 19.00
3 29 =(0.5*20+0.5*19) 19.50
4 38 =(0.5*29+0.5*19.5) 24.25
5 26 =(0.5*38+0.5*24.25) 31.13
6 28 =(0.5*26+0.5*31.125) 28.56
7 37 =(0.5*28+0.5*28.5625) 28.28
8 24 =(0.5*37+0.5*28.28125) 32.64
9 25 =(0.5*24+0.5*32.640625) 28.32
10 30 =(0.5*25+0.5*28.3203125) 26.66